OpenStudy (anonymous):

Find the solution set for y = 2x + 1, given the replacement set {(-2, -3), (-1, -1), (0, 2), (1, 3), (2, 6)}

OpenStudy (anonymous):

Test which values are true by plugging in the first value of each ordered pair in the place of x: For the first one: 2(-2)+1=-3, so solution set contains -2. For the second one 2(-1)+1=-1, so the solution set contains -1 For the third one 2(0)+1=1 (not 2), so the solution set does not contain 0. Just need to do the same for the last two.

OpenStudy (anonymous):

man i really don't get it!

jimthompson5910 (jim_thompson5910):

What this is asking is this: which ordered pair of the given set satisfies the given equation? For instance, does the ordered pair (-2,-3) satisfy the given equation y = 2x+1 The ordered pair (-2,-3) means that x = -2 and y = -3. So plug x = -2 and y = -3 into the equation above to get -3 = 2(-2)+1. Now evaluate the right side to get 2(-2)+1 = -4+1 = -3 So this means that the equation -3 = 2(-2)+1 then becomes -3 = -3. So the two values of x and y work. Therefore, they satisfy the equation. To determine if the rest of the ordered pairs work, you need to do the same as shown above. Hopefully that's enough to get you started. If not, then let me know.

OpenStudy (anonymous):

so it would be {(-2, -3), (-1,-1), (0, 2), (1, 3)} ?

jimthompson5910 (jim_thompson5910):

You're close, but (0,2) is NOT a solution because plugging in x = 0 and y = 2 means that you have y = 2x+1 2 = 2(0)+1 2 = 0+1 2 = 1 which is NOT true. So (0,2) is NOT a solution. Otherwise, you got it perfectly. Good job.

OpenStudy (anonymous):

thankyou so much!

OpenStudy (purple_pink):

@MCLover1477